OpenStudy (anonymous):

could someone help me with some algebra 2 problems?

OpenStudy (md152727):

yes

OpenStudy (md152727):

please post your question

OpenStudy (anonymous):

@MD152727

OpenStudy (anonymous):

First. Factor \[3x^2-17x-6\] by find factors of \[3(-6)\] whose sum is -17. This gives you: \[\frac{4x^2-32x+48}{x(3x+1)-6(3x+1)}\] Factor common terms:\[\frac{4x^2-32x+48}{(3x+1)(x-6)}\] Now factor out the greatest common divisor, at the numernator: \[\frac{4(x^2-8x+12)}{(3x+1)(x-6)}\] Now factor it:\[\frac{4(x-2)(x-6)}{(3x+1)(x-6)}\] Now cancel the terms: \[\frac{4(x-2)}{3x+1}\]

OpenStudy (anonymous):

Hope it makes sense @coltonbasham

OpenStudy (anonymous):

I am still confused on what my answer would be i see the work and understand that but i dont know

OpenStudy (anonymous):

@Tommynaiter

OpenStudy (anonymous):

You answer is \[\frac{4(x-2)}{3x+1}\] Which is one of your options :) ?

OpenStudy (anonymous):

@Tommynaiter could you help wit another one

OpenStudy (anonymous):

@Hero could you help me

OpenStudy (anonymous):

do you still need help?

OpenStudy (cutiekawdfi):

4